OpenStudy (anonymous):

Divide: 50a^11/ 4a^16

OpenStudy (asnaseer):

are you familiar with the rules to do with indexes?

OpenStudy (asnaseer):

e.g.:\[a^n/a^m=a^{n-m}\]

OpenStudy (anonymous):

\[{x^a \over x^b} = x^{a-b}\]

OpenStudy (asnaseer):

if not, then you might find this problem a bit more difficult

OpenStudy (asnaseer):

you can think of it as:\[\frac{50a^{11}}{4a^{16}}=\frac{50*a*a*a...*a}{4*a*a*a...*a}\] where the numerator has 11 multiples of "a" and the denominator has 16 multiples of "a"

OpenStudy (asnaseer):

so all you need to do is to cancel the common factors in this fraction. can you do that?

OpenStudy (anonymous):

So 50 * a^11/4a^16 = (50/4) * a^11/a^16= = 12,5 * a-5 or 12,5/a^5 If you prefer, 50/4*a^5 to

OpenStudy (anonymous):

alright got it